OpenStudy (anonymous):

find derivative of y=6root of x^5

OpenStudy (anonymous):

\[\sqrt[6]{x^5}\]

OpenStudy (abb0t):

\[\sqrt[6]{x^5} = x^{\frac{ 5 }{ 6 }}\]

OpenStudy (johnweldon1993):

as @abb0t has shown...those 2 expressions are equal....now how do you find the derivative of THAT?

OpenStudy (abb0t):

I don't know.

OpenStudy (johnweldon1993):

hint...think of the power rule

OpenStudy (anonymous):

it would be 5x/6?

OpenStudy (anonymous):

because you would get \[\frac{ 5 }{ 6 }X ^{\frac{ -1 }{ 6 }}\]

OpenStudy (anonymous):

which then to get rid of the negative you put the denominator of the negative integer as the denominator of the problem right?

OpenStudy (johnweldon1993):

well think of \[x ^{-1/2}\] That would become \[1/x ^{1/2}\]

OpenStudy (anonymous):

so it would be \[\frac{ 5 }{ 6x ^{1/6} }\]

OpenStudy (anonymous):

is that right?

OpenStudy (johnweldon1993):

That IS what I get....not 100% sure if it's corrct but I'll check again

OpenStudy (johnweldon1993):

yeah i keep getting that...anyone else have an opinion here?

OpenStudy (mathstudent55):

It's correct.

OpenStudy (abb0t):

i don't know.

OpenStudy (mathstudent55):